OpenStudy (anonymous):

Need Help Please! Not just the answer! Find the quadratic equation with roots -1+4i and -1-4i

jimthompson5910 (jim_thompson5910):

If the roots are -1+4i or -1-4i then x = -1+4i or x = -1-4i

jimthompson5910 (jim_thompson5910):

the goal is to get everything in each equation to the left side. This is so you are left with 0 on the right side

OpenStudy (anonymous):

so 0=-x-1-4i?

jimthompson5910 (jim_thompson5910):

x = -1 + 4i x+1 = -1 + 4i + 1 ... Add x to both sides. x+1 = 4i

jimthompson5910 (jim_thompson5910):

recall that i = sqrt(-1) to get rid of that 'i', you square both sides (to undo the square root)

OpenStudy (anonymous):

how do you sqrt x+1?

jimthompson5910 (jim_thompson5910):

so let's do so... x+1 = 4i (x+1)^2 = (4i)^2 (x+1)^2 = 4^2*i^2 (x+1)^2 = 16i^2 (x+1)^2 = 16(-1) (x+1)^2 = -16 now we can move the last bit to the other side (x+1)^2+16 = -16+16 ... Add 16 to both sides. (x+1)^2 + 16 = 0

jimthompson5910 (jim_thompson5910):

you square (x+1) NOT square root (x+1)

OpenStudy (anonymous):

Oh! Okay I got you.

jimthompson5910 (jim_thompson5910):

so you see how I got (x+1)^2 + 16 = 0 ?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

ok let's do the same for x = -1 - 4i

jimthompson5910 (jim_thompson5910):

you'll see that the steps are nearly identical

OpenStudy (anonymous):

Okay, and then do you do the quadratic formula ?

jimthompson5910 (jim_thompson5910):

x = -1-4i x+1 = -4i (x+1)^2 = (-4i)^2 (x+1)^2 = 16i^2 (x+1)^2 = 16(-1) (x+1)^2 = -16 (x+1)^2 + 16 = 0

jimthompson5910 (jim_thompson5910):

so as you can see, the two roots lead to (x+1)^2 + 16 = 0

jimthompson5910 (jim_thompson5910):

you use the quadratic formula to find the roots if you know the quadratic

jimthompson5910 (jim_thompson5910):

In this case, we don't know the quadratic but we know the roots

jimthompson5910 (jim_thompson5910):

anyways, the last step is to expand and simplify (x+1)^2 + 16

OpenStudy (anonymous):

So x^2 + 2x +17?

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

now if you used the quadratic formula for x^2 + 2x +17, you'll find that the roots are -1+4i and -1-4i

jimthompson5910 (jim_thompson5910):

so this is like thinking backwards

OpenStudy (anonymous):

Okay I get it now . Thank you so much for your help

jimthompson5910 (jim_thompson5910):

you're welcome